Calculus I, Fall 2013 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 08

1. Section 4.4: problems 77, 85, 105, 11.

2. Section 4.5: problems 21, 51, 57, 65, 73, 80, 81(b), 84(c), 86.

3. Section 4.6: problems 12, 27, 48, 62, 67(a).